What is the base and height of a triangle if the area is 98 meters squared and the height is 3/4 of the base?
Answer by venugopalramana(3286)

What is the base and height of a triangle if the area is 98 meters squared and the height is 3/4 of the base?
BASE = B
HEIGHT = (3/4)*BASE=3B/4
AREA = (1/2)BH=(1/2)*B(3B/4)=3B^2/8=98
B^2=8*98/3
B= SQRT(8*98/3)= 16.17
H=3*16.17/4 = 12.13